OpenStudy (anonymous):

(6.0*10^5)*(3.0*10^4)= ?

OpenStudy (anonymous):

answers are: 18*10^20 18*10^9 1.8*10^20 1.8*10^10 in scientific notation

OpenStudy (anonymous):

6*3 * 10^5*10^4 = 18 * 10 ^ 5+4 = 18 * 10^9 = 1.8 * 10^10

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

thanks